# SQL-Factory: A Multi-Agent Framework for High-Quality and Large-Scale SQL Generation

### v4

@@ Abstract.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
@@ 1. Introduction
Insufficient query diversity. A lack of diversity means that the generated SQL queries tend to follow similar structures, use the same types of operators, or involve limited combinations of joined tables. Template-based approaches such as TPC-DS (Nambiar and Poess, 2006) and Join Order Benchmark (JOB) (Leis et al., 2018) typically rely on fixed syntactic forms, which restrict their ability to produce varied query types. Learning-based methods such as TreeGAN (Liu et al., 2018) and LearnedSQLGen (Zhang et al., 2022) exhibit strong tendencies to overfit to frequent query skeletons in training corpora, producing workloads clustered around recurrent table subsets and operator chains. As a result, these methods fail to capture the rich variety of queries that appear in real-world workloads. Improving query diversity would help cover more edge cases and make trained models more robust in practical database environments.
@@
To address these limitations, we introduce SQL-Factory, a multi-agent framework designed for high-quality and large-scale SQL query generation. SQL-Factory leverages three collaborative agent teams: the Management Team, the Generation Team, and the Expansion Team, each responsible for a distinct aspect of the generation process. The Generation Team explores structurally diverse and expressive SQL patterns using a powerful large-scale language model, ensuring sufficient variety across query forms and structures. The Expansion Team complements this by scaling the workload through a lightweight and cost-effective local model, which generate semantically coherent variants based on representative seed queries, significantly reducing reliance on expensive models. Finally, the Management Team evaluates query quality and schema-level distribution in real time, guiding scheduling decisions to ensure that structurally complex or underrepresented schema components are adequately covered. This division of responsibilities enables SQL-Factory to achieve a promising trade-off between query diversity, generation efficiency, and schema-level balance.
In summary, the key contributions of this paper can be outlined as follows:
@@ 2.1. Data-Driven Intelligent Database
Another important line of work focuses on converting natural language to SQL (NL2SQL). A wide range of methods (Pourreza and Rafiei, 2023; Wang et al., 2025b; Gao et al., 2024; Zhang et al., 2024; Xu et al., 2025; Dong et al., 2023) train neural models on aligned natural language and SQL query pairs, enabling users to express complex information needs in plain language and receive executable queries as output.
Beyond these areas, research in this domain has also extended to cardinality estimation (Negi et al., 2023; Li et al., 2023b; Wang et al., 2023), index selection (Yu et al., 2024; Wang et al., 2025a), and query clustering. These applications benefit from the pattern recognition and generalization abilities of modern neural models.
@@ 2.2. SQL Generation
Early methods for SQL generation were primarily rule-based or template-driven. For example, industry-standard benchmarks such as TPC-DS (Nambiar and Poess, 2006), TPC-H (Poess and Floyd, 2000) and Join Order Benchmark (JOB) (Leis et al., 2018) construct queries from pre-defined templates, offering a controllable and interpretable means of workload construction. However, due to the limited number of templates, these methods often struggle to cover diverse query patterns and exhibit limited query randomness or structural variability. Random generation methods (Seltenreich, 2016; Slutz, 1998; Abdul Khalek and Khurshid, 2010) attempt to increase diversity by randomly walking through SQL grammar rules or syntax trees. These methods increase lexical diversity but lack semantic grounding, often producing unexecutable or meaningless queries that fail to reflect realistic workload patterns.
@@ 2.3. LLM-based Multi-Agent System
Building upon this foundation, recent research has extended LLM capabilities into collaborative multi-agent systems, where multiple specialized agents coordinate to complete tasks through dynamic communication and role-based reasoning. Multi-agent systems distribute cognitive responsibilities across distinct agents, each tailored for specific subgoals or expertise areas. This paradigm fosters more interpretable, scalable, and resilient behavior, particularly in settings that demand iterative planning, task delegation, and adaptive feedback handling. Multi-agent LLM systems have shown promise in domains such as software engineering (Qian et al., 2024; Chen et al., 2023), where agents representing roles like developer, tester, and reviewer to build and refine software artifacts. Similarly, in scientific experimentation (Zheng et al., 2023) and medical diagnostics (Tang et al., 2024), these systems enable complex reasoning chains by leveraging role specialization and interactive deliberation among agents.
In the context of databases, LLM-based multi-agent approaches have recently emerged as a powerful solution for tasks requiring high modularity and semantic precision. For instance, in Text-to-SQL translation, multiple agents can collaboratively handle subtasks such as schema linking, query generation, and final SQL refinement, leading to more robust and accurate SQL outputs (Wang et al., 2025b; Xie et al., 2024). In the realm of database anomaly diagnosis, systems like D-Bot (Zhou et al., 2024) leverage several large language models to automatically analyze and identify root causes of anomalies by extracting knowledge from diagnostic documents. Additionally, ROMAS (Huang et al., 2024) introduces a role-based multi-agent framework for database monitoring and planning, featuring agents assigned to specific roles (planner, monitor, and worker) collaboratively facilitate dynamic task allocation, self-monitoring, and adaptive re-planning in complex data environments. These examples collectively demonstrate the potential of multi-agent architectures to enhance reasoning modularity, improve decision accuracy, and integrate system-level feedback more effectively. Despite these advances, the application of LLM-based multi-agent systems to large-scale SQL generation remains underexplored.
@@ 3. System Overview
Expansion Team. The Expansion Team addresses the cost-efficiency challenge by scaling up the number of queries using lightweight generation. It includes the Seed Selection Agent and the Expansion Agent. The Seed Selection Agent identifies high-quality queries that involve under-covered tables, and the Expansion Agent uses them as seed queries to synthesize new variants via controlled augmentation. The Expansion Team concentrates on scaling promising patterns to achieve high query volume with minimal computational overhead, enabling the system to achieve both depth and breadth in SQL generation.
@@ 4.1.1. Management Agent
Specifically, the system begins in the exploration phase, where the Generation Team synthesizes queries for structurally complex tables or those with few synthesized queries. After a fixed number of generation cycles, control shifts to the Expansion Team to exploit learned patterns. During this exploitation phase, the Management Agent continuously monitors evaluation information from the Critical Agent. If query redundancy increases significantly or executability declines, it triggers a transition back to exploration. This alternating control mechanism allows SQL-Factory to incrementally grow a dataset that remains both diverse and scalable, while preventing the generation process from collapsing into repetitive or overly common query patterns.
@@ 4.1.2. Critical Agent
We introduce this hybrid metric to address the inherent limitations of embedding-based cosine similarity, which primarily captures surface-level semantic proximity. While effective for general clustering tasks, cosine similarity often fails to capture deeper structural or logical differences between SQL queries that have similar tokens but differ significantly in their execution semantics or syntax trees. To produce a more robust and semantically grounded similarity measure, we integrate multiple perspectives into a unified scoring framework. An illustration of the hybrid similarity computation is provided in Figure 3.
Formally, the similarity between a candidate query $q$ and a historical query $q^{\prime}\in\text{SQL Pool}$ is computed as:
@@
where $\alpha$, $\beta$, and $\gamma$ are the weights assigned to the three components. Empirically, we set $\alpha=0.6$, $\beta=0.3$, and $\gamma=0.1$ to balance between surface-level lexical overlap, deeper structural and semantic alignment. Each component here captures a different aspect of query similarity:
@@
    where $\text{EditDist}(\cdot)$ expresses the edit distance between AST trees, and ||$\cdot$|| ⋅ | represents the total number of nodes in the AST. This reflects the structural overlap between query patterns, including shared clauses and logical operators.
@@
    (5) $\text{Sim}_{\text{Emb}}(q,q^{\prime})=\frac{\text{Emb}(q)\cdot\text{Emb}(q^{\prime})}{||\text{Emb}(q)||\ ||\text{Emb}(q^{\prime})||},$
    where $\text{Emb}(q)$ denotes the vector representation of query $q$ encoded by the BERT encoder, and |||||$\cdot$||| | ⋅ | | represents the norm of the vector.
@@ 4.2.1. Table Selection Agent
(8) $\mathcal{T}^{\prime}=\text{Agent}_{\text{T}}(\mathcal{S}),\quad\text{where}\ \mathcal{T}^{\prime}\subset\mathcal{T}$
By integrating schema-level complexity and query statistics, the Table Selection Agent enables informed decision-making that balances exploration of new structural patterns with targeted improvement of under-sampled schema regions. This modular design also allows the system to dynamically adapt its generation focus based on the evolving distribution of queries in the SQL Pool.
@@ 4.2.2. Generation Agent
The agent is implemented using a powerful, instruction-tuned large language model, which is prompted with both schema information for $\mathcal{T}^{\prime}$ and high-level generation instructions, as shown in Figure 4. Rather than treating each table independently, the model considers the joint schema of all selected tables and is encouraged to synthesize complex queries that include multi-table joins, nested subqueries, aggregations, and window functions. Besides, to guide the model effectively, the schema information included in the prompt is carefully constructed to ensure both syntactic validity and semantic meaningfulness. For each selected table, we encode not only its column names and data types, but also relevant content-aware hints to reduce the likelihood of generating vacuous or logically invalid queries. Specifically, for columns of enumerated types (e.g., status, category), we provide up to five representative values from the domain to help the model infer plausible filter conditions or join constraints. For numeric or date type columns (e.g., int, float, date), we include its minimum and maximum values in the database, offering statistical context for range-based predicates (e.g., price ¿ 100). For free-text or string columns, we randomly sample up to five values as textual examples, assisting the model in generate meaningful LIKE conditions or equality conditions. These content-aware augmentations are inserted inline with the schema description in a structured format, forming an enriched prompt that provides the model with both structural knowledge and semantic information. This design significantly improves the quality and contextual relevance of the generated queries, particularly for tables with sparse usage history or rich attribute semantics. Finally, the output of the Generation Agent are summarized as:
@@
The generation is carried out in an open-ended fashion, without referencing queries from the SQL Pool. To ensure structural richness, prompts are explicitly designed to avoid trivial query types such as single-table scans or simple projections. This setup allows the Generation Agent to push the boundaries of the query space and introduce SQL patterns that improve structural diversity.
All generated queries $\mathcal{Q}_{\text{gen}}$ are forwarded to the Critical Agent for executability validation. Only those that compile successfully are inserted into the SQL Pool. In this way, the Generation Agent supports horizontal dataset expansion, discovering new structural dimensions of the underlying database schema.
@@ 4.3.2. Expansion Agent
Specifically, this agent operates on the set of seed queries, which were selected based on their relevance to underrepresented tables. It uses a lightweight model to derive multiple variants based on these seed queries, employing techniques such as structural rewriting, substructure recombination. Besides, the agent is not limited to structural transformations but also introduces lexical and logical variations to ensure semantic diversity.
@@
After expansion, the resulting queries $\mathcal{Q}_{\text{exp}}$ are passed to the Critical Agent. In this phase, queries are not only checked for executability but also evaluated for semantic redundancy using retrieval-based similarity assessment. Only queries that meet the criteria for both executability and novelty are retained and added to the SQL Pool. Through this process, the Expansion Agent supports vertical dataset expansion. It strengthens the density of high-quality query patterns within the local neighborhood of schema space, ultimately improving the dataset’s expressiveness and utility in downstream tasks. While the Generation Agent explores new structural forms, the Expansion Agent intensifies local pattern regions, ensuring both diversity and volume.
@@ 5.1.2. Benchmarks
To evaluate the effectiveness and generality of SQL-Factory, we select four typical database benchmarks for synthesizing queries. (1) TPC-DS (Nambiar and Poess, 2006) is an industry-standard benchmark for decision support systems, featuring a complex schema with 24 tables and various star-join patterns. It serves as a rigorous testbed for generating analytical SQL queries. (2) IMDB (Maas et al., 2011) is a movie-related database, which includes 21 tables on movies and actors. Its real-world structure and moderate size make it suitable for generating realistic join-heavy queries. (3) SPIDER (Yu et al., 2018) is a large-scale text-to-SQL dataset consisting of several databases with highly diverse and normalized schemas. We use the SPIDER schema collection as a target set for multi-domain query generation. (4) BIRD (Li et al., 2023a) is a recent benchmark introduced to evaluate database-instruction reasoning in realistic and complex environments. We use BIRD both as a target for SQL generation and as a downstream testbed for Text-to-SQL evaluation, particularly to assess robustness under schema ambiguity. Across all benchmarks, we parse the schema into structured representations including table names, column types, and foreign key relations. These are used as inputs for SQL-Factory’s Generation and Expansion workflows.

### v5

@@ Abstract.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
@@ 1. Introduction
Insufficient query diversity. A lack of diversity means that the generated queries tend to follow similar structures, use the same operators, or involve limited combinations of joined tables. Template-based approaches such as TPC-DS (Nambiar and Poess, 2006) and Join Order Benchmark (JOB) (Leis et al., 2018) typically rely on fixed syntactic forms, which restrict their ability to produce varied query types. Learning-based methods such as TreeGAN (Liu et al., 2018) and LearnedSQLGen (Zhang et al., 2022) exhibit strong tendencies to overfit to frequent query skeletons in training corpora, producing workloads clustered around recurrent table subsets and operator chains. As a result, these methods fail to capture the rich variety of queries that appear in real-world workloads. Improving query diversity would help cover more edge cases and make trained models more robust in practical database environments.
@@
To address these limitations, we introduce SQL-Factory, a multi-agent framework designed for high-quality and large-scale SQL query generation. SQL-Factory leverages three collaborative agent teams: the Management Team, the Generation Team, and the Expansion Team, each responsible for a distinct aspect of the generation process. The Generation Team explores structurally diverse and expressive SQL patterns using a powerful large-scale language model, ensuring sufficient variety across query forms and structures. The Expansion Team complements this by scaling the workload through a lightweight and cost-effective local model, which generate semantically coherent variants based on representative seed queries, significantly reducing reliance on expensive models. Finally, the Management Team evaluates query quality and schema-level distribution in real time, guiding scheduling decisions to ensure that structurally complex or underrepresented schema components are adequately covered. This division of responsibilities enables SQL-Factory to achieve a promising trade-off between query diversity, generation efficiency, and schema-level balance. Our contributions are summarized as follows:
@@ 2.1. Data-Driven Intelligent Database
Another important line of work focuses on converting natural language to SQL (NL2SQL). A wide range of methods (Pourreza and Rafiei, 2023; Wang et al., 2025b; Gao et al., 2024; Zhang et al., 2024; Xu et al., 2025; Dong et al., 2023; Kobayashi et al., 2025) train neural models on aligned natural language and SQL query pairs, enabling users to express complex information needs in plain language and receive executable queries as output.
Query clustering is also an important task, which aims to group structurally or semantically similar SQL queries in order to enable log summarization, workload compression, and behavior profiling. Prior work has proposed a variety of approaches based on syntactic structures (Kul et al., 2018) and learned representations (Tang et al., 2022). These methods typically require access to a diverse and well-structured corpus of queries to achieve meaningful clustering outcomes. However, such corpora are often unavailable or incomplete in real-world scenarios.
Beyond these applications, research in this domain have also been applied to cardinality estimation (Negi et al., 2023; Li et al., 2023b; Wang et al., 2023) and index selection (Yu et al., 2024; Wang et al., 2025a), where models benefit from the ability to generalize across query patterns and data distributions. These advances reflect the broader trend of integrating deep learning methods into database system components.
@@ 2.2. SQL Generation
Early methods for SQL generation were primarily rule-based or template-driven. For example, industry-standard benchmarks such as TPC-DS (Nambiar and Poess, 2006), TPC-H (Poess and Floyd, 2000) and Join Order Benchmark (JOB) (Leis et al., 2018) construct queries from pre-defined templates, offering a controllable and interpretable means of workload construction. Tenet(Bussotti et al., 2023) similarly adopts a template-driven approach, but introduces seed annotations and query types to synthesize more semantically diverse SQL queries. However, due to the limited number of templates, these methods often struggle to cover diverse query patterns and exhibit limited query randomness or structural variability. Random generation methods (Seltenreich, 2016; Slutz, 1998; Abdul Khalek and Khurshid, 2010) attempt to increase diversity by randomly walking through SQL grammar rules or syntax trees. These methods increase lexical diversity but lack semantic grounding, often producing unexecutable or meaningless queries that fail to reflect realistic workload patterns.
@@ 2.3. LLM-based Multi-Agent System
Building upon this foundation, recent research has extended LLM capabilities into collaborative multi-agent systems, where multiple specialized agents coordinate to complete tasks through dynamic communication and role-based reasoning. Multi-agent systems distribute cognitive responsibilities across distinct agents, each tailored for specific subgoals or expertise areas. Multi-agent LLM systems have shown promise in domains such as software engineering (Qian et al., 2024; Chen et al., 2023b), where agents representing roles like developer, tester, and reviewer to build and refine software artifacts. Similarly, in scientific experimentation (Zheng et al., 2023) and medical diagnostics (Tang et al., 2024), these systems enable complex reasoning chains by leveraging role specialization and interactive deliberation among agents.
In the context of databases, LLM-based multi-agent approaches have recently emerged as a powerful solution for tasks requiring high modularity and semantic precision. For instance, in Text-to-SQL translation, multiple agents can collaboratively handle subtasks such as schema linking, query generation, and final SQL refinement, leading to more robust and accurate SQL outputs (Wang et al., 2025b; Xie et al., 2024). In the realm of database anomaly diagnosis, systems like D-Bot (Zhou et al., 2024) leverage several large language models to automatically analyze and identify root causes of anomalies by extracting knowledge from diagnostic documents. Additionally, ROMAS (Huang et al., 2024) introduces a role-based multi-agent framework for database monitoring and planning, featuring agents assigned to specific roles (planner, monitor, and worker) collaboratively facilitate dynamic task allocation, self-monitoring, and adaptive re-planning in complex data environments. These examples demonstrate the potential of multi-agent architectures to enhance reasoning modularity, improve decision accuracy, and integrate system-level feedback more effectively. Despite these advances, the application of LLM-based multi-agent systems to large-scale SQL generation remains underexplored.
@@ 3. System Overview
Expansion Team. The Expansion Team addresses the cost-efficiency challenge by scaling up the number of queries using lightweight generation. It includes the Seed Selection Agent and the Expansion Agent. The Seed Selection Agent identifies high-quality queries that involve under-covered tables from the SQL Pool, and the Expansion Agent uses them as seed queries to synthesize new variants via controlled augmentation. The Expansion Team concentrates on scaling promising patterns to achieve high query volume with minimal computational overhead, enabling the system to achieve both depth and breadth in SQL generation.
@@ 4.1.1. Management Agent
Specifically, the system begins in the exploration phase, where the Generation Team synthesizes queries for structurally complex tables or those with few synthesized queries. After a fixed number of generation cycles, control shifts to the Expansion Team to exploit learned patterns. During this exploitation phase, the Management Agent continuously monitors evaluation information from the Critical Agent. If query redundancy increases significantly or executability declines, it triggers a transition back to exploration. This alternating control mechanism allows SQL-Factory to incrementally grow a dataset that remains both diverse and scalable, while preventing the generation process from collapsing into repetitive or overly common query patterns. Compared to rule-based heuristics or reinforcement learning, our LLM-based scheduling leverages schema statistics and top-k query neighbors returned by the Critical Agent to assess diversity and redundancy. This enables the agent to make informed, context-sensitive decisions across benchmarks without the need for retraining or hand-crafted rules.
@@ 4.1.2. Critical Agent
We introduce this hybrid metric to address the inherent limitations of embedding-based cosine similarity, which primarily captures surface-level semantic proximity. While effective for general clustering tasks, cosine similarity often fails to capture deeper structural or logical differences between queries that have similar tokens but differ significantly in their execution semantics or syntax trees. To produce a more robust and semantically grounded similarity measure, we integrate multiple perspectives into a unified scoring framework. An illustration of the hybrid similarity computation is provided in Figure 3. Formally, the similarity between a candidate query $q$ and a historical query $q^{\prime}\in\text{SQL Pool}$ is computed as:
@@
where $\alpha$, $\beta$, and $\gamma$ are the weights assigned to the three components. Each component here captures a different aspect of query similarity:
@@
    where $\text{EditDist}(\cdot)$ expresses the edit distance between AST trees, and $|\cdot|$ represents the total number of nodes in the AST. This reflects the structural overlap between query patterns, including shared clauses and logical operators.
@@
[…]
    where $\text{Emb}(q)$ denotes the vector representation of $q$ encoded by the BERT, and $||\cdot||$ represents the norm of the vector.
Given the distinct statistical behavior and discriminative power of these metrics, we assign the three components unequal weights, with $\alpha=0.6$, $\beta=0.3$, and $\gamma=0.1$, to reflect their differing variance ranges and discriminative effectiveness. In particular, token-level similarity exhibits the widest score dispersion and is thus most effective for detecting fine-grained redundancy. AST-level and embedding-level signals, while complementary, exhibit narrower and more concentrated distributions, making them more suitable as auxiliary signals.
@@ 4.2.1. Table Selection Agent
(8) $\mathcal{T}^{\prime}=\text{Agent}_{\text{T}}(\mathcal{S}),\quad\text{where}\ \mathcal{T}^{\prime}\subset\mathcal{T}.$
This agent operates over the full schema, and the underlying LLM can implicitly consider factors such as the number of columns, referential degree, and depth of the referential tree when selecting tables. These schema-level indicators highlight structural importance and connectivity. At the same time, the agent also accounts for query statistics, giving higher priority to tables that are currently underrepresented in the SQL Pool. In this way, both structurally complex tables and those with relatively few generated queries are emphasized. By integrating schema-level complexity with query-level statistics, the Table Selection Agent allows informed decision-making that balances exploration of new structural patterns with targeted improvement of under-sampled schema regions. This design enables dynamic adaptation as the SQL Pool evolves, ensuring diversity as well as representativeness in the generated queries.
@@ 4.2.2. Generation Agent
The agent is implemented using a powerful, instruction-tuned large language model, which is prompted with both schema information for $\mathcal{T}^{\prime}$ and high-level generation instructions, as shown in Figure 4. Rather than treating each table independently, the model considers the joint schema of all selected tables and is encouraged to synthesize complex queries that include multi-table joins, nested subqueries, aggregations, and window functions. Besides, to guide the model effectively, the schema information included in the prompt is carefully constructed to ensure both syntactic validity and semantic meaningfulness. In particular, we also incorporate explicit foreign key relationships, allowing the model to generate meaningful join conditions that align with actual schema constraints. For each selected table, we encode not only its column names and data types, but also relevant content-aware hints to reduce the likelihood of generating vacuous or logically invalid queries. Specifically, for columns of enumerated types (e.g., status, category), we provide up to five representative values from the domain to help the model infer plausible filter conditions or join constraints. For numeric or date type columns (e.g., int, float, date), we include its minimum and maximum values in the database, offering statistical context for range-based predicates (e.g., price ¿ 100). For free-text or string columns, we randomly sample up to five values as textual examples, assisting the model in generate meaningful LIKE conditions or equality conditions. These content-aware augmentations are inserted inline with the schema description in a structured format, forming an enriched prompt that provides the model with both structural knowledge and semantic information. This design improves contextual relevance and guides the model away from implausible or extreme filter conditions, enhancing overall semantic plausibility. This is particularly beneficial for tables with sparse usage history or complex attributes. Finally, the output of the Generation Agent are summarized as:
@@
The generation is carried out in an open-ended fashion, without referencing queries from the SQL Pool. To ensure structural richness, prompts are explicitly designed to avoid trivial query types such as single-table scans or simple projections. This setup allows the Generation Agent to push the boundaries of the query space and introduce SQL patterns that improve structural diversity. All generated queries $\mathcal{Q}_{\text{gen}}$ are forwarded to the Critical Agent for executability validation. Only those that compile successfully are inserted into the SQL Pool. In this way, the Generation Agent supports horizontal dataset expansion, discovering new structural dimensions of the underlying database schema.
@@ 4.3.2. Expansion Agent
Specifically, this agent operates on the set of seed queries, which were selected based on their relevance to underrepresented tables. It uses a lightweight model to derive multiple variants from these seed queries, employing techniques such as structural rewriting (e.g., rewriting subqueries as CTEs), substructure recombination (e.g., merging partial queries sharing common tables), lexical variation (e.g., alias renaming or clause reordering), and logical variation (e.g., modifying filter thresholds or switching conjunctions). These transformations are guided implicitly by the prompt rather than fixed rules, enabling flexible and semantically rich local exploration.
@@
After expansion, the resulting queries $\mathcal{Q}_{\text{exp}}$ are passed to the Critical Agent. In this phase, queries are checked for executability and evaluated for semantic redundancy using retrieval-based similarity assessment. Only queries satisfying both executability and novelty criteria are retained in the SQL Pool. Through this process, the Expansion Agent achieves vertical dataset growth by increasing the density of high-quality query patterns within the local schema space, enhancing the dataset’s expressiveness and utility for downstream tasks. While the Generation Agent explores new structural forms, the Expansion Agent intensifies local pattern regions, ensuring both diversity and volume.
@@ 5.1.2. Benchmarks
To evaluate the effectiveness and generality of SQL-Factory, we select four typical database benchmarks for synthesizing queries. (1) TPC-DS (Nambiar and Poess, 2006) is an industry-standard benchmark for decision support systems, featuring a complex schema with 24 tables and various star-join patterns. (2) IMDB (Maas et al., 2011) is a movie-related database, which includes 21 tables on movies and actors. Its real-world structure and moderate size make it suitable for generating realistic join-heavy queries. (3) SPIDER (Yu et al., 2018) is a large-scale text-to-SQL dataset consisting of several databases with highly diverse and normalized schemas. (4) BIRD (Li et al., 2023a) is a recent benchmark introduced to evaluate database-instruction reasoning in realistic and complex environments. Across all benchmarks, we parse the schema into structured representations including table names, column types, and foreign key relations. These are used as inputs for SQL-Factory’s Generation and Expansion workflows.